# Computational Design and Fabrication of Modular Robots with Untethered Control

### v1

@@ Acknowledgments
The authors express gratitude to Magali Lorion for assisting in the initial fabrication of LCEs, Yuhan Zhang for setting up the initial algorithm for the computational design tool, Pengbin Tang for providing the code for simulating discrete elastic rods, the Imaging and Optics Facility at ISTA for assisting with the spectrometry measurements, and the MIBA machine shop at ISTA for their support in manufacturing various devices.
Appendix A Supplementary Information
@@
The maximum contraction, the required laser actuation, and the actuation strength of the LCEs rods are function of their cross-sectional dimensions. LCEs were fabricated with varying widths and thicknesses, and through empirical evaluation, 5mm5\ $m$ width and 2mm2\ $m$ thickness were obtained as optimal dimensions. These dimension exhibited the minimal fixity by the LCE rods, good contraction capability ($33\%$) and provided reliable actuation behavior. A standard LCE rod is synthesized in a mold of dimension $60\times 5\times 2\ mm^{3}$. The synthesized LCEs were mechanically stretched to 90mm90\ $m$ and subsequently UV-cured, retaining this extended length at room temperature. Upon heating to 120°C they contract back to 60mm60\ $m$, achieving a contraction ratio of $33\%$. Each LCE rod of length 90mm90\ $m$ after fabrication weighs approximately 0.5g0.5\ g and is capable of lifting 100g100\ g of mass when heated to $120\degree C$. This corresponds to the LCE muscle generating a force of roughly 1N1\ N while maintaining its contracted length. Each muscle is cut to a specific length, according to the corresponding edge length in the robot’s skeletal graph.
@@
LCE synthesis is performed under experimental conditions, which can lead to fabrication discrepancies observable in the behavior of the LCE muscles. They may not exhibit the same response when actuated with a laser. This inconsistency is primarily due to the non-uniform distribution of the IR820820 dye, which results in clumped and uneven heating under infrared laser activation. To address this, a closed-loop control system is implemented to monitor each bone and its corresponding muscles, adjusting the actuation profile based on the current state. Furthermore, the laser system has a finite capacity and restricts the number of LCEs it can simultaneously heat. We explain this in the subsequent section. Moreover, since the LCEs are glued onto the bones, a few millimeters of each LCE become nonresponsive. This imposes a minimum LCE length of 30mm30\ $m$, which is required for reliable deformation. LCEs fabricated below this length fail to consistently achieve the desired maximum contraction of $33\%$. The minimum LCE length of 30mm30\ $m$ is used as a fabrication constraint (Eq. 2) in the design tool.
@@
A key constraint in any robotic design is the number of actuators that can be actuated concurrently. This limitation is dictated by both the physical characteristics of the LCE muscles and the performance constraints of the untethered laser actuation system. To estimate the number of actuators that can be admitted by our system, we conduct the following experiment. We actuate a LCE muscle of length 30mm30\ $m$ with the laser setup. We generate different experimental configurations ($Ei$) by varying laser speed and the number of repetitions. For each experimental configuration, we measure the operation time ($Oi_{t}$) of the laser to activate the LCE muscle and the cooling time ($Ci_{t}$) for the LCE muscle before it comes back to its pre-heated configuration. (see Fig. 3 (d)). During the cooling period, the laser is available to actuate other LCE muscles. We estimate the number of LCE muscles that can be actuated in parallel under configuration $Ei$ as
$N(Ei)=\frac{Ci_{t}}{Oi_{t}}$
We then select the configuration that maximizes $N(Ei)$ and obtained the maximum value as $30$ actuators. Therefore, under the current setup, we can simultaneously heat 30 LCE muscles, each of length 30mm30\ mm to a complete contraction of $33\%$, giving us a total contracted length of $30\times 30\ mm\times 33\%=300\ mm$. Most robotic designs, however, do not require full contraction of each edge, but partial contraction is often sufficient to achieve the desired shape or motion. Thus, for an edge $e_{i}$, with an initial length $l(e_{i})$ and a desired target length $d(e_{i})$, we obtain the actuation budget as the following constraint:
@@
This inequality serves as the feasibility constraint for any robot design and is incorporated into the design tool (see Eq. 2).
@@
We use the skeletal graph structure to generate the CAD file for the robot. The online CAD software Onshape [56] was used for this task. Each node is modeled using its vertex degree and the angle between the neighboring edges that are connected to this vertex. The position of the LCE slots is optimized using a Mathematica script [57] to ensure minimum spatial usage by the nodes, as they themselves do not contribute to actuation. To model the edges, we use connectors with a circular cross-section to connect the two nodes. If the edge has to be shrunk, an additional block part is generated in the middle to glue the LCEs. Finally, all the individual nodes and bones are generated using our script and are assembled in the OnShape assembly tool following the skeletal graph’s connectivity. The final CAD model is then ready to be used for 3D printing or simulation.
@@
We used an Ultimaker S7, a Fused Deposition Modeling 3D printer to manufacture the skeletons. The 3D printer was used with Ultimaker’s TPU-95A and Ultimaker’s PLA (Pearl White) materials. Ultimaker’s Natural PVA was used as support material. 3D prints made with TPU material are elastic, allowing for both shrinkage and preferential bending of the bones. However, skeletons made from TPU are prone to sagging under gravity. Bones printed with PLA are rigid and can only perform preferential bending, but they compensate with structural stiffness that enables them to maintain geometric constraints regardless of gravitational effects. Moreover, TPU is tricky to print in combination with the support material provided by Ultimaker making the overall print process quite cumbersome, whereas printing with PLA is relatively hassle free. Additionally, the frog-bot robot, was printed using the Bambu Lab X1C commercial robot with their PLA Basic material, which provides a similar mechanical response to Ultimaker PLA Pearl White, but the 3D-printed skeletons were more durable.
@@
We employed the AprilTag 16h5 dictionary, which can handle 30 unique IDs, enabling us to track multiple markers simultaneously. These ArUco markers are tracked by six RGB cameras (MC031CG-SY-UB, Ximea). To prevent imaging interference during laser operation, each RGB camera is equipped with an ND filter (NENIR260B, THORLABS) that blocks the infrared radiation emitted by the laser. Furthermore, to avoid burning the markers with the laser, the marker is printed in red on a white background. We used a standard corner detection and point triangulation functions from the OpenCV library to obtain the 3D location of the center and the plane normal of the markers. With our stereo-camera tracking we can determine the 3D coordinates of our marker with upto 2mm2\ $m$ accuracy. The current robotic designs use fewer than 30 nodes in the skeletal graph. If more nodes are required, AprilTag 36h11 dictionary can be used.
Figure 7: a) High lumen light for proper illumination, Infrared Laser and the 2-axis galvanoscanner setup. b) Teensy-4 and the power supply unit for the laser. c) The laser calibration and tracking setup comprising of 6 RGB, 1 monochrome and 1 thermographic camera. morph-bot visualized with monochrome camera (d), thermographic camera (e) and RGB camera (g). Control window in (f).
@@
Figure 8: Application of our computational design tool (Sec. 5: Computational Design for Shape Morphing Structures) to obtain highly-tessellated shape morphing structures. Our algorithm takes in multiple input shapes as input (in light-green) and generates a common triangulation for the skeletal graph of the flat-fabricated structure and the approximated target shapes (in dark-green). We use the [27] framework to 3D print five flat-fabricated structures that are visually the same but each is encoded with different shrinkage ratio per edge. These different flat-fabricated structures are passively actuated by heating in hot water and deform to their target configuration (in gray) closely approximating the input shapes.
Algorithm Details: Computational Design for Shape Morphing Structures
The design tool take as input $n$ different three dimensional manifold meshes[$\Omega^{1}_{0}(V^{1}_{0},F^{1}_{0})$, $\Omega^{2}_{0}(V^{2}_{0},F^{2}_{0})$$\Omega^{n}_{0}(V^{n}_{0},F^{n}_{0})$] and provide $n$ corresponding approximated meshes in 3D $\Omega^{1\star}(V^{1\star},T^{\star})$, $\Omega^{2\star}(V^{2\star},T^{\star})$$\Omega^{n\star}(V^{n\star},T^{\star})$] and a flat two dimensional mesh $\Gamma^{\star}(V^{flat\star},T^{\star})$ as output, all having the same connectivity $T^{\star}$. Mesh $\Omega^{i}$ represents the current state of the $i^{th}$ mesh, and $\Gamma$ represents the current flat mesh during the optimization process. Below we describe various energy functions that we use during the optimization process.
The map distortion energy minimizes the conformal distortion between approximate meshes $\Omega^{i}(V^{i},T)$ and $\Omega^{j}(V^{j},T)$ and ensures that the map remains as conformal as possible. This is important as the length ratio between the edges is related to the edge scale factor $^{\prime}u^{\prime}$ introduced in the Cherrier formulation as noted in [51]. We use these length ratios to compute the target muscle lengths of the robots. $E_{map}(\Omega^{i},\Omega^{j})$ between meshes $\Omega^{i}$ and $\Omega^{j}$ is defined between triangle $ti\in\Omega^{i}$ and its corresponding triangle $tj\in\Omega^{j}$ with their Jacobian $\mathcal{J}$ mapping $ti\rightarrow tj$, and the area $A_{t^{i}}$ and $A_{t^{j}}$ to produce the following:
@@
The pair wise combination between meshes are added up to yield the final conformal map distortion energy as:
$E_{map}(\Omega^{i})=\sum_{j\neq i}^{n}E_{map}(\Omega^{i},\Omega^{j})$
The mesh quality energy $E_{mesh}$ is used to prevent the triangles of the meshes from degenerating. It is defined between the triangles $ti\in\Omega^{i}$ and standard triangle $t^{eq}$ with their Jacobian as $J^{eq}(t^{i}\rightarrow t^{eq})$ and their area being $A_{t^{i}}$ and $A_{t^{eq}}$. To ensure that all the triangles $t\in T$ have similar size, the length of standard triangle is chosen to be the current average length of the edges in $\Omega^{i}$. $\epsilon^{eq}$ (= $0.3$) is used to tune the angle penalty to allow for variation of the standard triangle to be between equilateral and right angle triangles.
$E_{mesh\_area}(\Omega^{i})$ $=\left(\det(J^{eq})-1)\right)^{2}A_{t^{eq}}+\left(\det(J^{eq^{-1}})-1\right)^{2}A_{t}$
@@
$E_{mesh}(\Omega^{i})$ $=E_{mesh\_area}(\Omega^{i})+E_{mesh\_angle}(\Omega^{i})$
Since we work with disk-topology contrary, as opposed to spherical embeddings used by [52], the boundary of the mesh $\Omega^{i}$ needs to be convex to ensure that the embeddings of different meshes are bijective to each other. This is realized by keeping the boundary vertices $\mathcal{B}^{i}$ of the mesh $\Omega^{i}$ to lie on a unit disk. The boundary energy $E_{boundary}$ is defined as:
@@
The Bijective energy $E_{bij}$ ensures that the map between the approximated 3D mesh to its own embedding remains bijective. For a mesh $\Omega^{i}$, let $SA_{t}$ be the signed area of the triangle $t$ in its 2D embedding, the bijective energy is defined as:
@@
$E_{bij}(\Omega^{i})$ $=\sum_{t\in\Omega^{i}}E_{barrier}(t)$
The vertices of $\Omega^{i}$ lie on the surface of $\Omega^{i}_{0}$ by construction but the reverse statement is not true. The surface approximation energy measures this disparity. The vertices $v$ of $\Omega^{i}_{0}$ with normalized voronoi area defined as $area(v)$ are projected on $\Omega^{i}$ using barycentric coordinates ($\alpha,\beta,\gamma$) in the triangle (a,b,c) computed from their 2D embeddings using the inverse map $f^{-1}$ to obtain $\bar{v}$. The surface approximation energy $E_{approx}$ is defined as:
$\bar{v}$ $=\alpha f^{-1}(\mathbf{a})+\beta f^{-1}(\mathbf{b})+\gamma f^{-1}(\mathbf{c})$
@@
We formulate the shrinkage constraint (max 33% shrinkage, $\hat{r}=0.33$), imposed by our laser actuation setup between edge $ei\in\Gamma$ and $ej\in\Omega^{i}$, $r(ei,ej)=\frac{ei-ej}{ei}$ as a soft energy function where $E_{shrink}(r)$ is defined with $r_{0}=\epsilon_{3}$ and $r_{1}=\hat{r}-\epsilon_{3}$ as contraction constraints and $\epsilon_{0}$ and $\epsilon_{1}$ are used to soften the energy formulation. $\epsilon_{2}$ is initialized with $0.25$ and is gradually decreased to zero as the optimization progresses and $\epsilon_{3}=0.005$ to allow for small error buffer.
$\epsilon_{0}$ $=\max(0,\ \epsilon_{3}-\min_{r}(r))+\epsilon_{2}$
$\epsilon_{1}$ $=\max(0,\max_{r}(r)-\hat{r}+\epsilon_{3})+\epsilon_{2}$
$E_{shrink}(r)$ $=\begin{cases}-log(1+\frac{r-r_{0}}{\epsilon_{0}})+\frac{r-r_{0}}{\epsilon_{0}},&\text{if }-\epsilon_{0}+r_{0}<r<r_{0}\\ 0,&\text{if }r_{0}\leq r<r_{1}\\ -log(1-\frac{r-r_{1}}{\epsilon_{1}})-\frac{r-r_{1}}{\epsilon_{1}},&\text{if }r_{1}\leq r<r_{1}+\epsilon_{1}\\ \infty,&otherwise\end{cases}$
$E_{shrink-const}(\Gamma,\Omega^{i})$ $=\sum_{e\in T}E_{shrink}\bigg{(}r(\Gamma(e),\Omega^{i}(e))\bigg{)}$
Finally, the fabrication energy is defined as
$E_{fab}(\Gamma)=E_{bij}(\Gamma)$ $+w_{fab-map}\frac{1}{n}.\sum_{i=1}^{n}E_{map}(\Gamma,\Omega^{i})$
$+\frac{w_{fab-shrink}}{n}\sum_{i=1}^{n}E_{shrink-const}(\Gamma,\Omega^{i})$
@@
We weigh the different energy functions to obtain the combined energy function ($\mathcal{E}$) as:
$\mathcal{E}=\frac{1}{n}.\sum_{i=1}^{n}\Big{[}$ $\quad w_{bij}.E_{bij}(\Omega^{i})\ +w_{map}.E_{map}(\Omega^{i},\Gamma)\ +$
$\quad w_{mesh}.E_{mesh}(\Omega^{i},\Gamma)\ +w_{approx}.E_{approx}(\Omega^{i}_{0},\Omega^{i})\ +$
$\quad w_{boundary}.E_{boundary}(\Omega^{i})\Big{]}\ +w_{fab}.E_{fab}(\Gamma)\quad$
@@
Optimization: We first normalize the input meshes $\Omega_{0}^{i}$ to have unit surface area. We then obtain the embeddings of the normalized input meshes as the conformal parametrization and constrain them to lie on a unit disk using the method proposed by [51]. We initialize the flat mesh $\Gamma_{0}$ as the embedding of the first input mesh and use the triangulation of $\Gamma_{0}$ as the common triangulation $T_{0}$. We optimize the energy function $\mathcal{E}$ (3) defined above using the discrete continuous method as described in the algorithm 1 by [52] that we modify to satisfy our fabrication constraint (Eq. 2). The discrete step of the algorithm applies remeshing operation to the triangulation $T$ with the help of edge-flips, edge-splits and edge-collapse to reduce the energy function. While the continuous step of the algorithm, optimizes the vertex positions of the embeddings of the approximated meshes $\Omega^{i}$ and the flat mesh $\Gamma$. The optimization is performed in two different stages: a coarse stage and a fabrication stage. Each stage is run for 100 iterations, and the tuned weights are reported in Table 2. The pseudocode in Algorithm 1 gives an overview of this process.
Table 2: Parameters for different stages of the optimization

### v2

@@ Acknowledgments
The authors express gratitude to Magali Lorion for assisting in the initial fabrication of LCEs, Pengbin Tang for providing the code for simulating discrete elastic rods, the Imaging and Optics Facility at ISTA for assisting with the spectrometry measurements, and the MIBA machine shop at ISTA for their support in manufacturing various devices.
Funding:
This project was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant Agreement No. 715767 -– MATERIALIZABLE).
Author contributions:
M.B. conducted the primary research for this study, including the synthesis of LCE material, the characterization of the 3D-printed bone structures, and the characterization of the laser heating setup with LCE muscles. M.B. also designed, fabricated, and executed experiments for the various robotic systems presented; automated the CAD design pipeline; and developed the design and simulation tools described in the paper. M.B. was responsible for drafting the original manuscript and preparing the figures. Additionally, M.B. contributed to the building and calibration of the laser setup. T.H. designed, built, and calibrated the laser actuation hardware and software and implemented the marker tracking system. T.H. also assisted in the design of the laser-based experiments. M.S. advanced the research on LCE muscles by developing the IR-absorbing LCE materials used in the study, and conducted experiments for the mechanical characterization of the bones and muscles. M.B and M.S. designed the building block of the robots and the experimental environments for the locomotive robots. Y.Z. contributed to the computational design tool for shape morphign structures. M.P. contributed to the conceptualization of the design and simulation tools. B.B. conceived the study. C.D. served in an advisory role and helped with conceptualization. D.I. and B.B. supervised the project. All authors contributed to project discussions and the editing of the manuscript.
Competing interests:
The authors declare that they have no competing interests.
Data and materials availability:
The code for numerical simulations has been deposited in the GitHub repository: https://github.com/manas-avi/CDFMRUC.
Supplementary Information
@@
The maximum contraction, the required laser actuation, and the actuation strength of the LCEs rods are functions of their cross-sectional dimensions. LCEs were fabricated with varying widths and thicknesses, and through empirical evaluation, 5mm5\ $m$ width and 2mm2\ $m$ thickness were obtained as optimal dimensions. These dimensions exhibited the minimal fixity by the LCE rods, good contraction capability ($33\%$) and provided reliable actuation behavior. A standard LCE rod is synthesized in a mold of dimension $60\times 5\times 2\ mm^{3}$. The synthesized LCEs were mechanically stretched to 90mm90\ $m$ and subsequently UV-cured, retaining this extended length at room temperature. Upon heating to 120°C they contract back to 60mm60\ $m$, achieving a contraction ratio of $33\%$. Each LCE rod of length 90mm90\ $m$ after fabrication weighs approximately 0.5g0.5\ g and is capable of lifting 100g100\ g of mass when heated to $120\degree C$. This corresponds to the LCE muscle generating a force of roughly 1N1\ N while maintaining its contracted length. Each muscle is cut to a specific length, according to the corresponding edge length in the robot’s skeletal graph.
@@
LCE synthesis is performed under experimental conditions, which can lead to fabrication discrepancies observable in the behavior of the LCE muscles. They may not exhibit the same response when actuated with a laser. This inconsistency is primarily due to the non-uniform distribution of the IR820820 dye, which results in clumped and uneven heating under infrared laser activation. To address this, a closed-loop control system is implemented to monitor each bone and its corresponding muscles, adjusting the actuation profile based on the current state. Furthermore, the laser system has a finite capacity and restricts the number of LCEs it can simultaneously heat. We explain this in the subsequent section. Moreover, since the LCEs are glued onto the bones, a few millimeters of each LCE become nonresponsive. This imposes a minimum LCE length of 30mm30\ $m$, which is required for reliable deformation. LCEs fabricated below this length fail to consistently achieve the desired maximum contraction of $33\%$. The minimum LCE length of 30mm30\ $m$ is used as a fabrication constraint (Equation 2) in the design tool.
@@
A key constraint in any robotic design is the number of actuators that can be actuated concurrently. This limitation in our system is dictated by both the physical characteristics of the LCE muscles and the performance constraints of the untethered laser actuation system. To estimate the number of actuators that can be admitted by our system, we conduct the following experiment. We actuate a LCE muscle of length 30mm30\ $m$ with the laser setup. We generate different experimental configurations ($Ei$) by varying laser speed and the number of repetitions. For each experimental configuration, we measure the operation time ($Oi_{t}$) of the laser to activate the LCE muscle and the cooling time ($Ci_{t}$) for the LCE muscle before it comes back to its pre-heated configuration. (see Figure 3 (D)). During the cooling period, the laser is available to actuate other LCE muscles. We estimate the number of LCE muscles that can be actuated in parallel under configuration $Ei$ as
$N(Ei)=\frac{Ci_{t}}{Oi_{t}}.$
We then select the configuration that maximizes $N(Ei)$ and obtain the maximum value as $30$ actuators. Therefore, under the current setup, we can simultaneously heat 30 LCE muscles of length 30mm30\ $m$ each, with (maximum) contraction rate of $33\%$, giving us a total contracted length of $30\times 30\ mm\times 33\%=300\ mm$. Most robotic designs, however, do not require full contraction of each edge, but partial contraction is often sufficient to achieve the desired shape or motion. Thus, for an edge $e_{i}$, with an initial length $l(e_{i})$ and a desired target length $d(e_{i})$, we obtain the actuation budget as the following constraint:
@@
This inequality serves as the feasibility constraint for any robot design and is incorporated into the design tool (see Equation 2).
@@
We use the skeletal graph structure to generate the CAD file for the robot. The online CAD software Onshape  (?) was used for this task. Each node is modeled using its vertex degree and the angle between the neighboring edges that are connected to this vertex. The position of the LCE slots is optimized using a Mathematica script  (?) to ensure minimum spatial usage by the nodes, since they do not contribute to actuation. To model the edges, we use connectors with a circular cross-section to connect the two nodes. If the edge has to be shrunk, an additional block part is generated in the middle to glue the LCEs. Finally, all the individual nodes and bones are generated using our script and are assembled in the OnShape assembly tool following the skeletal graph’s connectivity. The final CAD model is then ready to be used for 3D printing or simulation.
@@
We used an Ultimaker S7, a Fused Deposition Modeling 3D printer to manufacture the skeletons. The 3D printer was used with Ultimaker’s TPU-95A and Ultimaker’s PLA (Pearl White) materials. Ultimaker’s Natural PVA was used as support material. 3D prints made with TPU material are elastic, allowing for both shrinkage and preferential bending of the bones. However, skeletons made from TPU are prone to sagging under gravity. Bones printed with PLA are rigid and can only perform preferential bending, but their structural stiffness enables them to maintain geometric constraints regardless of gravitational effects. Moreover, TPU is tricky to print in combination with the support material, making the overall print process quite cumbersome, whereas printing with PLA is relatively hassle free. Additionally, the frog-bot robot, was printed using the Bambu Lab X1C commercial printer with their PLA Basic material, which provides a similar mechanical response to Ultimaker PLA Pearl White, but the 3D-printed skeletons were more durable.
@@
We employed the AprilTag 16h5 dictionary, which can handle 30 unique IDs, enabling us to track multiple markers simultaneously. These ArUco markers are tracked by six RGB cameras (MC031CG-SY-UB, Ximea). To prevent imaging interference during laser operation, each RGB camera is equipped with an ND filter (NENIR260B, THORLABS) that blocks the infrared radiation emitted by the laser. Furthermore, to avoid burning the markers with the laser, the marker is printed in red on a white background. We used a standard corner detection and point triangulation functions from the OpenCV library to obtain the 3D location of the center and the plane normal of the markers. With our stereo-camera tracking, we can determine the 3D coordinates of our marker with up to 2mm2\ $m$ accuracy. The current robotic designs use fewer than 30 nodes in the skeletal graph. If more nodes are required, AprilTag 36h11 dictionary can be used.
Figure S1: (A) High lumen light for proper illumination, Infrared Laser and the 2-axis galvanoscanner setup. (B) Teensy-4 and the power supply unit for the laser. (C) The laser calibration and tracking setup comprising of 6 RGB, 1 monochrome and 1 thermographic camera. morph-bot visualized with monochrome camera (D), thermographic camera (E) and RGB camera (G). Control window in (F).
Figure S2: Application of our computational design tool (Computational Design Tool for Shape Morphing Materials and Methods: Computational Design for Shape Morphing Structures) to obtain highly-tessellated shape morphing structures. Our algorithm takes in multiple input shapes as input (in light-green) and generates a common triangulation for the skeletal graph of the flat-fabricated structure and the approximated target shapes (in dark-green). We use the  (?) framework to 3D print five flat-fabricated structures that are visually the same but each is encoded with different shrinkage ratio per edge. These different flat-fabricated structures are passively actuated by heating in hot water and deform to their target configuration (in gray) closely approximating the input shapes.
@@
Algorithm Details of Computational Design for Shape Morphing Structures
The design tool takes as input $N$ different three-dimensional manifold meshes $\{\Omega^{1}_{0}(V^{1}_{0},F^{1}_{0})$, $\Omega^{2}_{0}(V^{2}_{0},F^{2}_{0})$$\Omega^{n}_{0}(V^{n}_{0},F^{n}_{0})\}$ and output $N$ corresponding approximated meshes in 3D $\{\Omega^{1\star}(V^{1\star},T^{\star})$, $\Omega^{2\star}(V^{2\star},T^{\star})$$\Omega^{n\star}(V^{n\star},T^{\star})\}$ and a two-dimensional flat mesh $\Gamma^{\star}(V^{flat\star},T^{\star})$, all having the same connectivity $T^{\star}$. Mesh $\Omega^{i}$ represents the current approximated mesh of the input mesh $\Omega^{i}_{0}$, and $\Gamma$ represents the current flat mesh during the optimization process. Below we describe the details and rationale behind different energy functions, initialization and optimization process used by our design tool.
Initialization: We normalize the input meshes $\Omega_{0}^{i}$ to have unit surface area, then obtain embeddings of the normalized input meshes as the conformal parametrization map ($f$) and constrain them to lie on a unit disk using the method proposed by  (?). The inverse of this map $f^{-1}$ is used to lift vertices in the embedding space to the input mesh using the barycentric coordinates. We initialize the triangulation $T$ with two triangles and use this triangulation to remesh the original embeddings and initialize the 2D mesh $\Gamma$. The vertices of the remeshed embeddings and $\Gamma$ are chosen as points on the unit circle that maximize the area of the triangles. Although we define the energies below with respect to the approximated meshes $\Omega^{i}$}, the actual optimization variables are the remeshed embeddings. The approximated surfaces are obtained by lifting these embeddings through the inverse map $f^{-1}$.
Energy functions: Map Distortion Energy: $E_{map}(\Omega^{i},\Omega^{j})$ minimizes the conformal distortion between two approximate meshes $\Omega^{i}(V^{i},T)$ and $\Omega^{j}(V^{j},T)$ and ensures that the map remains as conformal as possible. A discrete conformal map is a piecewise linear map between triangulated surfaces that locally preserves angles by scaling edge lengths. Thus, preserving conformal map helps ensure that the two approximated meshes can be morphed into each other by simply contracting edges, which is the mechanism by which we actuate our robots. For a triangle $t\in T$, $J$ is defined as the Jacobian that maps the triangle $t^{i}\in\Omega^{i}$ to $t^{j}\in\Omega^{j}$ with $A_{t^{i}}$ and $A_{t^{j}}$ as their areas. Thus, the map distortion energy is formulated as:
@@
Mesh Quality Energy: $E_{mesh}$ is the weighted sum of different energy terms that help in preserving the quality of approximated meshes $\Omega^{i}$ and is defined as:
$E_{mesh}(\Omega^{i})=w_{tri}E_{tri}(\Omega^{i})+w_{bij}E_{bij}(\Omega^{i})+w_{boundary}E_{bounary}(\Omega^{i})+w_{approx}E_{approx}(\Omega^{i}).$
Triangle Quality Energy: $E_{tri}$ preserves the triangle quality of the approximating triangulations by penalizing the deviation of each triangle $ti$ from a target equilateral triangle $t^{eq}$. We define Jacobian $J^{eqi}(t^{i}\rightarrow t^{eq})$ and their area as $A_{t^{i}}$ and $A_{t^{eq}}$. Length of $t^{eq}$ is obtained by specifying the surface approximation parameter ($\epsilon_{approx}$) that adaptively computes the length of $t^{eq}$ based on the curvature of the mesh  (?). $\epsilon^{eq}$ (= $0.3$) is used to tune the angle penalty to allow for variation of the standard triangle to be between equilateral and right angle triangles. This gives us the following formulation:
$E_{tri\_area}(\Omega^{i})$ $=\sum_{t\in\Omega^{i}}\left(\det(J^{eq})-1)\right)^{2}A_{t^{eq}}+\left(\det(J^{eq^{-1}})-1\right)^{2}A_{t}\text{,}$
@@
$E_{tri}(\Omega^{i})$ $=E_{tri\_area}(\Omega^{i})+E_{tri\_angle}(\Omega^{i})\text{.}$
Bijective energy: $E_{bij}$ ensures that the map between the approximated 3D mesh to its remeshed embedding remains bijective by preventing the triangles from degenerating. For a mesh $\Omega^{i}$, let $SA_{t}$ be the signed area of the triangle $t$ in its 2D embedding, the bijective energy is defined as:
@@
$E_{bij}(\Omega^{i})$ $=\sum_{t\in\Omega^{i}}E_{barrier}(t)\text{.}$
Boundary Energy: $E_{boundary}$ controls the boundary of the remeshed embedding of mesh $\Omega^{i}$. Since we work with disk-topology, as opposed to spherical embeddings used by  (?), all vertices of the remeshed embedding need to stay within the unit circle for the map between remeshed embedding and approximated mesh to remain bijective. Furthermore, the remeshed embedding need to cover as much area of the unit circle to provide better approximation of the input surface. Both of these objectives are maintained by keeping the boundary vertices $\mathcal{B}^{i}$ of the remeshed embedding of $\Omega^{i}$ to lie on a unit disk. Thus giving the boundary energy as:
@@
Approximation energy: The vertices of $\Omega^{i}$ lie on the surface of $\Omega^{i}_{0}$ by construction but the reverse statement is not true. The $E_{approx}$ measures this disparity and helps improving the approximation of $\Omega^{i}$ towards $\Omega^{i}_{0}$. The vertices $v\in\Omega^{i}_{0}$ with normalized voronoi area $area(v)$ are projected on $\Omega^{i}$ from embedding space using barycentric coordinates to obtain $\bar{v}$. This gives us surface approximation energy formulation as:
[…]
Fabrication Energy: $E_{fab}$ consists of three energy terms designed to fulfill the fabrication constraints, i.e. bijective constrain, conformal constraint and shrinkage constraint:
$E_{fab}(\Gamma,\Omega^{i})=w_{bij}E_{bij}(\Gamma)+w_{fabmap}E_{map}(\Gamma,\Omega^{i})+w_{shrink}E_{shrink}(\Gamma,\Omega^{i}).$
@@
$E_{bij}$ and $E_{map}$ has the same formulation as before. $E_{bij}$ prevents the triangles in the 2D mesh $\Gamma$ from degenerating. $E_{map}$ optimize map from the 2D mesh $\Gamma$ to every approximated mesh $\Omega^{i}$ to be as conformal as possible mimicking the edge-contraction based deformation. We define the Shrinkage energy $E_{shrink}$ to enforce the maximum shrinkage constraint (33%) of the laser actuation setup. For every edge $e\in T$, the shrinkage rate is computed with the corresponding edges $e_{\gamma}\in\Gamma$ and $e_{i}\in\Omega^{i}$ as $r(e_{\gamma},e_{i})=\frac{len(e_{\gamma})-len(e_{i})}{len(e_{\gamma})}$. $E_{sh}(r)$ is defined as a soft energy function with $\hat{r}=0.33$ as the maximum shrinkage rate. $\epsilon_{0}$ and $\epsilon_{1}$ are used to soften the energy formulation. $\epsilon_{2}$ is initialized with $0.25$ and is gradually decreased to zero as the optimization progresses and $\epsilon_{3}$ with $0.005$ to allow for small error buffer.
$\epsilon_{0}$ $=\max(0,\ \epsilon_{3}-\min_{r}(r))+\epsilon_{2}\text{,}$
$\epsilon_{1}$ $=\max(0,\max_{r}(r)-\hat{r}+\epsilon_{3})+\epsilon_{2}\text{,}$
$r_{0}$ $=\epsilon_{3}\text{, }r_{1}=\hat{r}-\epsilon_{3}\text{,}$
$E_{sh}(r)$ $=\begin{cases}-log(1+\frac{r-r_{0}}{\epsilon_{0}})+\frac{r-r_{0}}{\epsilon_{0}}&\text{if }-\epsilon_{0}+r_{0}<r<r_{0},\\ 0&\text{if }r_{0}\leq r<r_{1},\\ -log(1-\frac{r-r_{1}}{\epsilon_{1}})-\frac{r-r_{1}}{\epsilon_{1}}&\text{if }r_{1}\leq r<r_{1}+\epsilon_{1},\\ \infty&\text{otherwise,}\end{cases}$
$E_{shrink}(\Gamma,\Omega^{i})$ $=\sum_{e\in T}E_{sh}\bigg{(}r(\Gamma(e),\Omega^{i}(e))\bigg{)}\text{.}$
@@
We weigh and normalize these energy functions and add them to obtain the combined energy function ($\mathcal{E}$) as defined in 1:
$\mathcal{E}=\frac{1}{{N\choose 2}}\sum_{ij}E_{map}(\Omega^{i},\Omega^{j})\ +\ \frac{1}{N}\sum_{i}E_{mesh}(\Omega^{i})\ +\ \frac{1}{N}\sum_{i}E_{fab}(\Gamma,\Omega^{i}).$
@@
Optimization: We obtain the desired approximated meshes $\Omega^{i\star}$ and $\Gamma^{\star}$ by optimizing the above defined energy function $\mathcal{E}$ (Algorithm Details of Computational Design for Shape Morphing Structures). We perform the optimization in two different stages. For each stage, the energy function $\mathcal{E}$ is optimized using the discrete continuous optimization method as described in the algorithm 1 by  (?) under our fabrication constraints (Equation 2). The discrete step of the algorithm applies remeshing operation including edge splits, edge collapses and edge flips to the triangulation $T$ in order to reduce the energy function. The continuous step of the algorithm then optimizes the vertex positions of the embeddings of the approximated meshes $\Omega^{i}$ and the flat mesh $\Gamma$. The weights of the first ‘map’ stage are chosen to optimize the approximation quality of $\Omega^{i\star}$ and the weights of the second ‘fab’ stage facilitates the optimization of $\Gamma^{\star}$. Each stage is run for 100 iterations, and the tuned weights are reported in Table S1. The pseudocode in 1 gives an overview of this process.
Table S1: Parameters for different stages of the optimization